Operations Report – April 11th

Crew 192 Operations Report 11Apr2018


Name of person filing report: Joe Dituri

Non-nominal systems: Structural – the main hab/EVA door.

Notes on non-nominal systems: The door swung open during the severe wind storm and knocked the top piece of the railing off. We secured the door as best we could and it swung open again. We secured it better and it’s holding for now.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 67% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas,

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 994 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133345.0

Toilet tank emptied: No. Last emptied April 9.

ATV’s Used: Red ATV, Blue ATV #2

Oil Added? No But checked and OK.

ATV Fuel Used: 5% Red, 15% Blue #2

# Hours the ATVs were Used today: 2

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Spirit rover used: Yes
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Crew is doing well, we had to stay on alert today for the strong winds. Will remain alert for continued winds tomorrow.

Summary of internet: 160MB at 1930.

Summary of suits and radios: Suits all charging, 2 one-piece suits used today, radios working fine.

Summary of Hab operations: Cleaned both decks, moved frig and stove to clean behind, used afternoon EVA to check up around habitat after strong winds, fixed cement step in front of rear airlock as it was not secure and almost fell apart.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: Suits inspected for safety and the Science Dome was cleaned and mopped.

Summary of RAM operations: The trailer water tank that we rolled into the RAM yesterday was examined further so we could devise the best way to clean it over the next few days.

Summary of health and safety issues: In an effort to sweep the upper deck and clean behind the refrigerator, we found a spilled box of matches on the floor near the wall directly between the fridge’s electrical system and propane stove. We removed the matches and swept the area. See Science Report for additional notes.

Questions, concerns, and requests to Mission Support: See Commander’s Report.

Copyright © The Mars Society. All rights reserved. | Main Site